OpenStudy (sleepyjess):

Help with a hyperbola question? Find an equation in standard form for the hyperbola with vertices at (0, +- 2) and foci at (0, +- 7).

OpenStudy (sleepyjess):

@ganeshie8

OpenStudy (sleepyjess):

I know that the equation will be \(\dfrac{y^2}{a^2} - \dfrac{x^2}{b^2} = 1\), but I'm not sure how to find a and b

ganeshie8 (ganeshie8):

Vertices = \((0,~\pm a\)) compare that with given vertices and find \(a\)

OpenStudy (sleepyjess):

So a would be 7!

OpenStudy (sleepyjess):

No, 2!

ganeshie8 (ganeshie8):

Yes \(a=2\), next use below and find \(c\) first Foci = \((0,\pm c)\)

OpenStudy (sleepyjess):

c would be 7 :)

ganeshie8 (ganeshie8):

use the raltion \(a^2+b^2=c^2\) and find \(b\)

OpenStudy (sleepyjess):

\(b^2\) would be 45

ganeshie8 (ganeshie8):

Yep! plug them in the standard form of equation of vertical hyperbola

OpenStudy (sleepyjess):

\(\dfrac{y^2}4 - \dfrac{x^2}{45}=1\)! "D

OpenStudy (sleepyjess):

* :D

ganeshie8 (ganeshie8):

looks good! you may double check with wolfram http://www.wolframalpha.com/input/?i=focus+y%5E2%2F4-x%5E2%2F45%3D1

OpenStudy (sleepyjess):

That was actually a lot easier than I thought it would be :)

ganeshie8 (ganeshie8):

it was easy because you remembered the standard form and formulas for vertices and foci :)

OpenStudy (sleepyjess):

:) I always overthink math